OpenStudy (anonymous):

let f(x)=-2x+4 and g(x)=-6-7. Find f(x)-g(x). PLEASE HELP!!

OpenStudy (wmj259):

Did you forget the x after the -6? for g(x)?

OpenStudy (anonymous):

ah yes, sorry D:

OpenStudy (anonymous):

f(x)-g(x)={-2x+4}-{-6x-7} =-2x+4+6x+7 (on removing brackets) =4x+13

OpenStudy (wmj259):